OpenStudy (anonymous):

7/19 = ?/114 Convert to higher terms.. I just dont get it

OpenStudy (anonymous):

you have to see what 19 should be multiplied with to get 114 then you multiply 7 by the same number to get the numerator for the RHS

OpenStudy (anonymous):

What must you multiply the 19 in the denominator of the first fraction to get 114?

OpenStudy (anonymous):

suppose the number is H, then 7 * H ? ------ = ------ 19 * H 114

OpenStudy (anonymous):

For converting equivalent fractions, you basically just multiply the fraction by some 'version' of 1: \[\frac{7}{19}\cdot \frac{2}{2} = \frac{14}{38}\] etc.

OpenStudy (anonymous):

So we are looking for the version of 1: \[\frac{7}{19} \cdot \frac{h}{h} = \frac{?}{114}\] If it makes the denominator 114 then what must h be? Then multiply that h by 7 to find the numerator.

OpenStudy (anonymous):

So like 19 times 6 = 114

OpenStudy (anonymous):

correct..

OpenStudy (anonymous):

so now calculate the numerator

OpenStudy (anonymous):

Right. So that means that h must be 6. Therefore we have: \[\frac{7}{19}\cdot \frac{6}{6} = \frac{??}{114}\] What is the ??

OpenStudy (anonymous):

got it????

OpenStudy (anonymous):

I got 42

OpenStudy (anonymous):

That's correct.

OpenStudy (anonymous):

absolutely correct !! Well Done !!! I suggest you do a few more such sums to become familiar with the concept and method....☺